Mrs oli plans to select 360 grade eleven students obtained from the population. The selected respondents has an average height of 170 cm and a sample standard deviation of 25 cm. What is the 95% confidence interval of all the student's heights?





1
Expert's answer
2022-05-30T15:24:21-0400

The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}},\bar{x}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(170-1.96\times\dfrac{25}{\sqrt{360}},170+1.96\times\dfrac{25}{\sqrt{360}})"

"=(167.418,172.582)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "167.418 < \mu < 172.582," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(167.418, 172.582)."